# LRSAA: Large-scale Remote Sensing Image Target Recognition and Automatic Annotation

### v1

@@ Abstract
This paper presents a method for object recognition and automatic labeling in large-area remote sensing images called LRSAA. The method integrates YOLOv11 and MobileNetV3-SSD object detection algorithms through ensemble learning to enhance model performance. Furthermore, it employs Poisson disk sampling segmentation techniques and the EIOU metric to optimize the training and inference processes of segmented images, followed by the integration of results. This approach not only reduces the demand for computational resources but also achieves a good balance between accuracy and speed. The source code for this project has been made publicly available on https://github.com/anaerovane/LRSAA.
@@ III Methods
Figure 2: Methods: Our methodology comprises four stages: Step A, B, C, and D. In Step A, we utilize Poisson disk sampling and segmentation techniques to partition the large-scale image into smaller segments. Step B involves training YOLOv11 and MobileNetV3-SSD models on these smaller images to detect objects. For Step C, the detection results from the smaller images are mapped back onto the original large-scale image to maintain spatial consistency. Finally, in Step D, we augment the original dataset with synthetic data generated through this process, allowing for re-training the models with an enriched dataset that includes proportional synthetic samples. This approach ensures continuous improvement and adaptation of the models to diverse scenarios.Figures 1 and 2 are sourced from Papers [6] and [11].
@@ IV-A Initial training
We utilized the XView dataset as our initial dataset. XView is a large-scale aerial image dataset containing over 1 million objects annotated with bounding boxes and class labels, designed for advancing computer vision tasks in satellite imagery analysis[21, xview2]. To align with the test image dimensions, we cropped the images from the XView dataset into smaller images of size 640×640 and remapped the label positions accordingly.

### v3

@@ Abstract
This paper introduces a novel method for object recognition and automatic labeling in large-area remote sensing images, called LRSAA. The proposed method integrates the YOLOv11 and MobileNetV3-SSD object detection algorithms through ensemble learning to enhance overall model performance. Additionally, it utilizes Poisson disk sampling segmentation techniques along with the EIOU metric to optimize both the training and inference processes of segmented images, culminating in an integrated results framework. This approach not only minimizes computational resource requirements but also achieves an effective balance between accuracy and processing speed. The source code for this project is publicly accessible at https://github.com/anaerovane/LRSAA.
@@ III Methods
Figure 2: Methods: Our methodology comprises four stages: Step A, B, C, and D. In Step A, we utilize Poisson disk sampling and segmentation techniques to partition the large-scale image into smaller segments. Step B involves training YOLOv11 and MobileNetV3-SSD models on these smaller images to detect objects. For Step C, the detection results from the smaller images are mapped back onto the original large-scale image to maintain spatial consistency. Finally, in Step D, we augment the original dataset with synthetic data generated through this process, allowing for re-training the models with an enriched dataset that includes proportional synthetic samples. This approach ensures continuous improvement and adaptation of the models to diverse scenarios.
@@ IV-A Initial training
We utilized the XView dataset as our initial dataset. XView is a large-scale aerial image dataset containing over 1 million objects annotated with bounding boxes and class labels, designed for advancing computer vision tasks in satellite imagery analysis[21]. To align with the test image dimensions, we cropped the images from the XView dataset into smaller images of size 640×640 and remapped the label positions accordingly.